Solutions of iron (III) chloride, FeCl3, are used in photoengraving and to make ink. This compound can be made by the following reaction:

2Fe + 3Cl sub 2, reaction arrow, 2FeCl sub 3.

How many grams of Fe are needed to combine with 4.5 moles of Cl2?
1
Expert's answer
2015-05-12T09:41:47-0400
From the reaction:
3 moles of Cl2 react with 2 moles of Fe,
So 4,5 moles of Cl2 will react with X moles of Fe.
X=4,5*2/3=3 moles of Fe
Molar mass of(Fe)=56 g/mole, so m(Fe)=3*56=168 grams
Answer: 168 grams